# Correspondences of the Third Kind: Camera Pose Estimation from Object Reflection

### v1

@@ Abstract
Computer vision has long relied on two kinds of correspondences: pixel correspondences in images and 3D correspondences on object surfaces. Is there another kind, and if there is, what can they do for us? In this paper, we introduce correspondences of the third kind we call reflection correspondences and show that they can help estimate camera pose by just looking at objects without relying on the background. Reflection correspondences are point correspondences in the reflected world, i.e., the scene reflected by the object surface. The object geometry and reflectance alters the scene geometrically and radiometrically, respectively, causing incorrect pixel correspondences. Geometry recovered from each image is also hampered by distortions, namely generalized bas-relief ambiguity, leading to erroneous 3D correspondences. We show that reflection correspondences can resolve the ambiguities arising from these distortions. We introduce a neural correspondence estimator and a RANSAC algorithm that fully leverages all three kinds of correspondences for robust and accurate joint camera pose and object shape estimation just from the object appearance. The method expands the horizon of numerous downstream tasks, including camera pose estimation for appearance modeling (e.g., NeRF) and motion estimation of reflective objects (e.g., cars on the road), to name a few, as it relieves the requirement of overlapping background.
@@ 1 Introduction
Recent neural shape-from-shading methods can recover the object geometry as surface normals in each view even under complex natural illumination [21]. As the illumination is unknown, the recovered surface normals, however, suffer from a fundamental ambiguity known as the generalized bas-relief ambiguity [1] between light source and surface geometry. That is, a rotated illumination and sheered surface can conspire to generate the exact same object appearance. As such, 3D correspondences established between the two surface normal maps recovered independently from the two images would only tell us the camera pose in the distorted space.
@@
How then can camera pose be estimated just from the object appearance? We take a hint from what we humans likely do: establish correspondences in the world reflected by the object surface. By leveraging the single-view reconstruction of surface normals, we can extract from each of the images a reflectance map. The reflectance map is the Gaussian sphere of the reflected radiance. It represents the surrounding environment modulated by surface reflectance as a spherical surface indexed by surface normals. We establish correspondences between the reflectance maps of the images. We refer to these as reflection correspondences.
We show that these reflection correspondences by themselves are not sufficient to directly compute the camera pose. They, however, resolve the ambiguity remaining in the point and 3D correspondences in presence of the bas-relief ambiguity. This means that, by using all three types of correspondences in two or more images of an object of arbitrary reflectance, even without any texture and with strong specularity, we can compute the relative camera poses just from the object appearances. This liberates many applications from seemingly benign yet practically extremely limiting requirement of overlapping static background or diffuse surface texture just to recover camera positions.
We formalize reflection correspondences and show how they should be combined with conventional correspondences (pixel correspondences and 3D correspondences) to obtain a quantitative estimate of the correct camera rotation. We also introduce a new neural feature extractor for establishing 3D and reflection correspondences robust to the inherent distortions primarily caused by the bas-relief ambiguity with effective data augmentation. Finally, we derive a RANSAC-based joint estimation framework for accurate joint camera pose and geometry reconstruction which alternates between the interdependent two quantities.
Experimental results on synthetic correspondences and images show that the reflection correspondences and the neural feature extractor as well as our iterative estimation framework are essential for accurate camera pose and geometry estimation. We believe reflection correspondences can play an important role in applications beyond camera pose and shape recovery from object appearance, including camera calibration [2] and object pose estimation [14] when classical correspondences are not sufficient.
@@ 2 Related Work
Structure-from-motion methods recover camera poses for multiple images by detecting and leveraging pixel correspondences [25, 16]. They typically detect such pixel correspondences by leveraging view-independent, salient features such as textures and jointly solve for camera poses and 3D locations of the surface points. Descriptors invariant to rotation and scale (e.g., SIFT [10]) and outlier correspondence detection with RANSAC [5] are often exploited for robust estimation. However, these methods are prone to fail on textureless, non-Lambertian objects (especially, shiny objects) as correspondences between surface points are extremely challenging to establish on these objects. Also, correspondences based on textures are usually sparse, and these methods require a large number (e.g., 50) of input images that have high visual overlaps for accurate estimation.
Neural image synthesis methods jointly estimate surface geometry and a surface light field (or its radiometric roots, i.e., reflectance and illumination) as neural representations with differentiable volumetric [11, 13, 24, 19, 9] or surface [22, 23] rendering. While most of these methods require multiple registered images as inputs, a few methods also handle non-registered multi-view images by optimizing the camera poses as additional parameters [22, 3]. They, however, still require good initial estimates of camera poses (e.g., SAMURAI [3] uses manually annotated coarse camera poses as inputs) or an extremely large number ($>$50) of input images as the joint optimization of all unknown parameters is extremely challenging and easily falls into local minima.
Pose estimation from specular reflections has also been studied [8, 15, 4, 12]. Lagger et al. [8], for instance, refine camera pose estimates using images and the 3D CAD model of a specular object. They recover view-dependent environment maps from the inputs and optimize the camera pose by minimizing discrepancies between them. These methods, however, assume known geometry [8, 15, 4, 12], known illumination [4], or a simple lighting model (i.e., directional or point source) [15, 12] which cannot be assumed in general scenes.
@@ 3 Method
In this section, we first detail the three types of correspondences we consider and the equations that can be derived from them. Figure 2 provides a visualization of these three types of correspondences. To avoid confusion, we refer to them by the following names:
  • “pixel correspondences”, for the “classical correspondences” which link corresponding locations in the two images on the object surface: the two matched points are the reprojections of the same physical 3D point in the two images.
  • “3D correspondences”, for the correspondences between corresponding normals in the two images on the object surface. Note that for each pixel correspondence, we also have a 3D correspondence if we know the normals at the matched image locations. The difference is that for the pixel correspondence, we exploit the pixel coordinates themselves, while in the case of the 3D correspondence, we exploit the normals.
  • “reflection correspondences”, which match image locations where light rays creating specular reflections come from the same direction in the two images. As shown in Fig. 2(c), we detect these correspondences from camera-view reflectance maps, i.e., view-dependent maps about the surrounding environment.
@@ 3.1 Pixel Correspondences
Let us assume that we can establish $N_{\mathrm{IM}}$ pixel correspondences $(u_{1}^{i},v_{1}^{i})\leftrightarrow(u_{2}^{i},v_{2}^{i})$. Under orthographic projection, each pixel correspondence gives us one constraint on the camera rotation between the two images [6]:
@@
where $\phi$ and $\theta$ are two of three Euler angles $(\theta,\phi,\eta)$ that correspond to the relative rotation $\mathrm{R_{21}}$ from the second view to the first one. Without loss of generality, we assume that these correspondences are shifted such that the average location (e.g., $\frac{1}{N_{\mathrm{IM}}}\sum_{i}u_{1}^{i}$) is zero. We use the (z-x-z) sequence for the Euler angles
@@
where $\mathrm{R_{z}}[\cdot]$ and $\mathrm{R_{x}}[\cdot]$ are rotations around z and x axes, respectively. Since Eq. 1 is independent of one of the unknown Euler angles ($\eta$), it is not possible to fully recover the rotation matrix from pixel correspondences when we have only two views.
@@
If we could estimate unambiguously the normal maps ${\bf N}_{1}$ and ${\bf N}_{2}$ for the images, correspondences $(u_{1}^{i},v_{1}^{i})\leftrightarrow(u_{2}^{i},v_{2}^{i})$ would be constrained by
@@
$\mu$ and $\nu$ can take any values, and $\lambda$ can take any positive value. Note that, while discussions regarding the bas-relief ambiguity (e.g., one in Belhumeur et al. [1]) usually assume directional lights, this holds true even for environmental illumination as we can see it as a set of directional lights.
@@
There is, however, an unresolvable ambiguity in its decomposition into the three unknown matrices. Most important, this ambiguity corresponds to one regarding the pixel correspondences, in other words, we cannot recover the relative camera pose even when combining pixel and 3D correspondences: for any $\eta$, there are corresponding GBR transformations $\mathrm{G_{1}}$ and $\mathrm{G_{2}}$ that are consistent with Eq. 6. We provide the proof in the supplementary material.
@@ 3.3 Reflection Correspondences
Let us assume we have correspondences between image locations $(u_{1}^{i},v_{1}^{i})\leftrightarrow(u_{2}^{i},v_{2}^{i})$ where light rays come from the same direction in the two images. In other words, correspondences of the surroundings reflected by the object surface. If we, for now, ignore the GBR ambiguity, these “reflection” correspondences give each a constraint of the form
@@
Each reflection correspondence thus gives us a new type of equations to estimate $\mathrm{R_{21}}$ but also $\mathrm{G_{1}}$ and $\mathrm{G_{2}}$, from which we can also get the object shape.
Figure 3: Given two-view images of a textureless, non-Lambertian object, we first recover surface normals and reflectance map for each view using a single-view geometry reconstruction method [21]. We establish 3D and reflection correspondences with a novel deep feature extraction network (Fig. 5) to compute the relative camera pose from them.
Detection of these reflection correspondences directly from images is, however, extremely challenging as surface reflection depends not only on the surrounding illumination but also on the surface geometry. Let us now assume we have a surface normal map for each view. Then we can avoid this problem by recovering camera-view reflectance maps [7]. The reflectance maps are view-dependent mappings from a surface normal to the surface radiance which are determined by the surface reflectance and the surrounding illumination environment
@@
where $\mathbf{\omega_{i}}$, $\mathbf{\omega_{o}}$, $\mathbf{n}$ are incident, viewing, and surface normal orientations in the local camera coordinate system, respectively. $\mathrm{R}_{k}$ is the camera pose (a rotation matrix) of $k$-th view, $L_{i}(\mathrm{v})$ is a mapping from an incident direction to a radiance of the the corresponding incident light, and $\psi(\mathbf{\omega_{i}},\mathbf{\omega_{o}},\mathbf{n})$ is the bidirectional reflectance distribution function (BRDF).
From the surface normal maps and the input images, i.e., pairs of surface normals and surface radiances, we can recover these reflectance maps and, as illustrated in Fig. 2(c), detect these reflection correspondences from the reflectance maps regardless of the object shape. Note that, as the reflectance maps are recovered using the surface normal maps, they also suffer from the bas-relief ambiguity. The relationship between the ground truth $E(\mathbf{n})$ and another possible solution $E(\mathbf{n})$ is
@@
3.4 Objective Function
From Eq. 1, (6), and (10), we derive the objective function
$f=f_{\text{IM}}+f_{\text{NM}}^{(12)}+f_{\text{NM}}^{(21)}+f_{\text{RM}}^{(12)}+f_{\text{RM}}^{(21)}\,.$ (13)
that enforces the equation in the least-squares sense. $f_{\text{IM}}$ enforces Eq. 1:
$f_{\text{IM}}=\sum_{i}^{N_{\text{IM}}}\left(t_{\phi}^{i}-t_{\theta}^{i}\right)^{2}\,,$ (14)
@@ 3.4 Objective Function
and $N_{\text{IM}}$ is the number of image correspondences.
@@
$f_{\text{NM}}^{(jk)}=\sum_{i}^{N_{\text{NM}}}\left\|{\bf N}_{j}[u_{j}^{i},v_{j}^{i}]-{\text{Norm}}(\mathrm{G}_{kj}{\bf N}_{k}[u_{k}^{i},v_{k}^{i}])\right\|^{2}\,,$ (17)
where Norm is the vector normalization operator and $N_{\text{NM}}$ the number of 3D correspondences.
@@
$f_{\text{RM}}^{(jk)}=\sum_{i}^{N_{\text{RM}}}\left\|{\bf N}_{j}[u_{j}^{i},v_{j}^{i}]-\Omega^{(jk)}({\bf N}_{k}[u_{k}^{i},v_{k}^{i}])\right\|^{2}\,,$ (18)
@@
Given 3 sets of $N_{\text{IM}}$ pixel correspondences, $N_{\text{NM}}$ 3D correspondences, and $N_{\text{RM}}$ reflection correspondences, we optimize $f$ in Eq. 13 for the three Euler angles $\theta$, $\phi$, and $\eta$, and the two sets of parameters for the GBR transformations for both images $\mu_{1}$, $\nu_{1}$, $\lambda_{1}$, $\mu_{2}$, $\nu_{2}$, $\lambda_{2}$, under the constraints that $\lambda_{1}$ and $\lambda_{2}$ are positive. We empirically find that naive nonlinear optimization with all the correspondences is slow and susceptible to local minima. We instead derive a two-step algorithm that first estimates the combined transformation $\mathrm{G_{21}}$ using pixel and 3D correspondences and then decomposes it using reflection correspondences. Nonlinear optimization with an off-the shelf solver [18] is used for the first step. In the second step, we discretize $\eta$ in 180 angles, solve for the other parameters analytically, and check if the obtained parameters are consistent with reflectance map correspondences. The supplementary material provides more details.
Intuition on the required numbers of correspondences.
In practice, the numbers of correspondences required for the camera pose estimation is important as we can use them for a robust estimation with RANSAC [5]. Although it is difficult to derive a rigorous proof about the number of required correspondences, we can build intuition about how many correspondences are required for each view pair. We first consider estimation from 3D correspondences in normal maps. Each correspondence gives us 1 constraint from the images and 2 from the normal maps—only 2 because, while the surface normals are 3-vectors, they are normalized. The number of unknown parameters except for the unresolvable parameter $\eta$ is 8. Thus the expected number of required 3D correspondences in normal maps is 3. One reflection correspondence is also required for resolving the remaining ambiguity about $\eta$. We indeed experimentally confirmed that three 3D and one reflection correspondences used together for direct optimization of all transformations ($\mathrm{G_{1}},\mathrm{G_{2}}$ and $R_{21}$) are sufficient. When using the sequential two-step algorithm, however, we empirically find in Sec. 4.1 that we need four 3D correspondences instead of three for the first stage of estimating the combined transformation $\mathrm{G_{21}}$. This is actually consistent with the total of 4 mixed 3D and reflection correspondences necessary for a direct optimization, likely suggesting that the least-square estimation of the combined transformation always needs at least a total of four correspondences of whatever form. Deriving a rigorous theoretical understanding of this would open new interesting research directions in 3D geometry.

### v2

@@ Abstract
Computer vision has long relied on two kinds of correspondences: pixel correspondences in images and 3D correspondences on object surfaces. Is there another kind, and if there is, what can they do for us? In this paper, we introduce correspondences of the third kind we call reflection correspondences and show that they can help estimate camera pose by just looking at objects without relying on the background. Reflection correspondences are point correspondences in the reflected world, i.e., the scene reflected by the object surface. The object geometry and reflectance alter the scene geometrically and radiometrically, respectively, causing incorrect pixel correspondences. Geometry recovered from each image is also hampered by distortions, namely generalized bas-relief ambiguity, leading to erroneous 3D correspondences. We show that reflection correspondences can resolve the ambiguities arising from these distortions. We introduce a neural correspondence estimator and a RANSAC algorithm that fully leverages all three kinds of correspondences for robust and accurate joint camera pose and object shape estimation just from the object appearance. The method expands the horizon of numerous downstream tasks, including camera pose estimation for appearance modeling (e.g., NeRF) and motion estimation of reflective objects (e.g., cars on the road), to name a few, as it relieves the requirement of overlapping background.
Keywords:
Camera Pose Reflection Bas-Relief Ambiguity
@@ 1 Introduction
Recent neural shape-from-shading methods can recover the object geometry as surface normals in each view even under complex natural illumination [36]. As the illumination is unknown, the recovered surface normals, however, suffer from a fundamental ambiguity known as the generalized bas-relief ambiguity [1] between light source and surface geometry. That is, a rotated illumination and sheared surface can conspire to generate the exact same object appearance. As such, 3D correspondences established between the two surface normal maps recovered independently from the two images would only tell us the camera pose in the distorted space.
@@
How then can the camera pose be estimated just from the object appearance? We take a hint from what we humans likely do: establish correspondences in the world reflected by the object surface. By leveraging the single-view reconstruction of surface normals, we can extract from each of the images a reflectance map. The reflectance map is the Gaussian sphere of the reflected radiance. It represents the surrounding environment modulated by surface reflectance as a spherical surface indexed by surface normals. We establish correspondences between the reflectance maps of the images. We refer to these as reflection correspondences.
We show that these reflection correspondences by themselves are not sufficient to directly compute the camera pose. They, however, resolve the ambiguity remaining in the pixel and 3D correspondences due to the bas-relief ambiguity. This means that, by using all three types of correspondences in two or more images of an object of arbitrary reflectance, even without any texture and with strong specularity, we can compute the relative camera poses just from the object appearance. This liberates many applications from the seemingly benign yet practically extremely limiting requirement of overlapping static background or diffuse surface texture just to recover camera positions.
We formalize reflection correspondences and show how they should be combined with conventional correspondences (pixel correspondences and 3D correspondences) to obtain a quantitative estimate of the correct camera rotation. We introduce a RANSAC-based, two-step algorithm which first exploits conventional correspondences and then resolves the ambiguity with reflection correspondences. We also introduce a new neural feature extractor for establishing 3D and reflection correspondences robust to the inherent distortions primarily caused by the bas-relief ambiguity with effective data augmentation. Finally, we derive a joint estimation framework for accurate joint camera pose and geometry reconstruction which alternates between the interdependent two quantities.
Experimental results on synthetic and real images show that the reflection correspondences and the neural feature extractor as well as our iterative estimation framework are essential for accurate camera pose and geometry estimation. We believe reflection correspondences can play an important role in applications beyond camera pose and shape recovery from object appearance, including camera calibration [2] and object pose estimation [26] when classical correspondences are not sufficient. All data and code can be found on our project web page.
@@ 2 Related Work
Structure-from-motion methods recover camera poses for multiple images by detecting and leveraging pixel correspondences [41, 30]. They typically detect such pixel correspondences by leveraging view-independent, salient features such as textures and jointly solve for camera poses and 3D locations of the surface points. Descriptors invariant to rotation and scale (e.g., SIFT [20]) and outlier correspondence detection with RANSAC [9] are often used for robust estimation. However, these methods are prone to fail on textureless, non-Lambertian objects (especially, shiny objects) as correspondences between surface points are extremely challenging to establish on these objects. Also, correspondences based on textures are usually sparse, and these methods require a large number (e.g., 50) of input images that have large visual overlaps for accurate estimation.
Neural image synthesis methods jointly estimate surface geometry and a surface light field (or its radiometric roots, i.e., reflectance and illumination) as neural representations with differentiable volumetric [22, 24, 40, 32, 19] or surface [37, 39] rendering. While most of these methods require multiple registered images as inputs, a few methods also handle non-registered multi-view images by optimizing the camera poses as additional parameters [37, 3]. They, however, still require good initial estimates of camera poses (e.g., SAMURAI [3] uses manually annotated coarse camera poses as inputs) or an extremely large number ($>$50) of input images as the joint optimization of all unknown parameters easily falls into local minima.
Pose estimation from specular reflection has also been studied [16, 29, 6, 23, 18, 11]. Lagger et al. [16], for instance, refine camera pose estimates using images and the 3D CAD model of a specular object. They recover view-dependent environment maps from the inputs and optimize the camera pose by minimizing discrepancies between them. These methods, however, assume known geometry [16, 29, 6, 23], known illumination [6, 18, 11], or a simple lighting model [29, 23] which cannot be assumed in general scenes.
@@ 3 Method
In this section, we first detail the three types of correspondences we consider and the equations that can be derived from them. Figure 2 provides a visualization of these three types of correspondences. To avoid confusion, we give distinct names to them.
  • “pixel correspondences” for the classical correspondences which link corresponding locations in the two images on the object surface: the two matched points are the reprojections of the same physical 3D point in the two images.
  • “3D correspondences” for the correspondences between corresponding normals in the two images on the object surface. Note that for each pixel correspondence, we also have a 3D correspondence if we know the normals at the matched image locations. The difference is that for the pixel correspondence, we exploit the pixel coordinates themselves, while in the case of the 3D correspondence, we exploit the normals.
  • “reflection correspondences” which match image locations where light rays mirror-reflected by the object surface come from the same direction in the two images. As shown in Fig. 2(c), we detect these correspondences from camera-view reflectance maps, i.e., view-dependent maps about the surrounding environment.
@@
Figure 3: As depicted in (3(a)) and (3(b)), due to the generalized bas-relief ambiguity in single-view surface normal recovery [1] and the fundamental difficulty in structure-from-motion [12], we cannot obtain a unique solution for the relative rotation from pixel and 3D correspondences when the cameras are orthographic (e.g., distant from the object). Reflection correspondences, i.e., correspondences regarding the incident directions for specular reflections, enable us to distinguish the correct relative rotation (3(c)) from the other possible solutions like (3(d)).
@@ 3.1 Pixel Correspondences
Let us assume that we can establish $N_{\mathcal{I}}$ pixel correspondences $(u_{1}^{i},v_{1}^{i})\leftrightarrow(u_{2}^{i},v_{2}^{i})$. Under orthographic projection, each pixel correspondence gives us one constraint on the camera rotation between the two images [12]:
@@
where $\phi$ and $\theta$ are two of the three Euler angles $(\theta,\phi,\eta)$ that correspond to the relative rotation $\mathrm{R_{21}}$ from the second view to the first one. Without loss of generality, we assume that these correspondences are translated such that the average location (e.g., $\frac{1}{N_{\mathcal{I}}}\sum_{i}u_{1}^{i}$) is zero. We use the (z-x-z) sequence for the Euler angles
@@
where $\mathrm{R_{z}}[\cdot]$ and $\mathrm{R_{x}}[\cdot]$ are rotations around z and x axes, respectively. As depicted in Figs. 3(a) and 3(b), since Eq. 1 is independent of one of the unknown Euler angles ($\eta$), the rotation matrix cannot be fully recovered from pixel correspondences of only two views.
@@
If we could estimate unambiguously the normal maps, ${\bf N}_{1}$ and ${\bf N}_{2}$, for the images, correspondences $(u_{1}^{i},v_{1}^{i})\leftrightarrow(u_{2}^{i},v_{2}^{i})$ would be constrained by
@@
The parameters $\mu$ and $\nu$ can take any values, and $\lambda$ can take any positive value. Note that, although discussions regarding the bas-relief ambiguity (e.g., one in Belhumeur et al. [1]) usually assume directional lights, this holds true even for environmental illumination as we can view it as a set of directional lights.
@@
There is, however, an unresolvable ambiguity in its decomposition into the three unknown matrices. Most important, this ambiguity corresponds to one regarding the pixel correspondences. In other words, we cannot recover the relative camera pose even when combining pixel and 3D correspondences: for any $\eta$, there are corresponding GBR transformations $\mathrm{G_{1}}$ and $\mathrm{G_{2}}$ that are consistent with Eq. 6. We provide the proof in the supplementary material.
@@ 3.3 Reflection Correspondences
Let us assume we have correspondences between image locations $(u_{1}^{i},v_{1}^{i})\leftrightarrow(u_{2}^{i},v_{2}^{i})$ where light rays come from the same direction in the two images. In other words, correspondences of the surroundings reflected by the object surface. If we, for now, ignore the GBR ambiguity, these “reflection correspondences” give each a constraint of the form
@@
As illustrated in Figs. 3(c) and 3(d), each reflection correspondence thus gives us a new type of equation to estimate $\mathrm{R_{21}}$ but also $\mathrm{G_{1}}$ and $\mathrm{G_{2}}$, from which we can also get the object shape.
Detecting these reflection correspondences directly from images is, however, extremely challenging as surface reflection depends not only on the surrounding illumination but also on the surface geometry. Let us now assume we have a surface normal map for each view. Then we can avoid this problem by recovering camera-view reflectance maps [13]. The reflectance maps are view-dependent mappings from a surface normal to the surface radiance which are determined by the surface reflectance and the surrounding illumination environment
@@
where $\mathbf{\omega_{i}}$, $\mathbf{\omega_{o}}$, and $\mathbf{n}$ are incident, viewing, and surface normal orientations in the local camera coordinate system, respectively. $\mathrm{R}_{k}$ is the camera pose (a rotation matrix) of $k$-th view, $L_{i}(\mathrm{v})$ is a mapping from an incident direction to a radiance of the corresponding incident light, and $\psi(\mathbf{\omega_{i}},\mathbf{\omega_{o}},\mathbf{n})$ is the bidirectional reflectance distribution function (BRDF).
From the surface normal maps and the input images, i.e., pairs of surface normals and surface radiances, we can recover these reflectance maps [27, 36] and, as illustrated in Fig. 2(c), detect these reflection correspondences from the reflectance maps regardless of the object shape. Note that, as the reflectance maps are recovered using the surface normal maps, they also suffer from the bas-relief ambiguity. The relationship between the ground truth $E(\mathbf{n})$ and another possible solution $E^{\prime}(\mathbf{n})$ is
@@
3.4 Relative Rotation from Correspondences
From Eq. 1, (6), and (10), we can derive an objective function
$f=f_{\mathcal{I}}+f_{\mathcal{N}}^{(12)}+f_{\mathcal{N}}^{(21)}+f_{\mathcal{R}}^{(12)}+f_{\mathcal{R}}^{(21)}\,.$ (13)
that enforces the equation in the least-squares sense. $f_{\mathcal{I}}$ enforces Eq. 1:
$f_{\mathcal{I}}=\frac{1}{{N_{\mathcal{I}}}}\sum_{i}^{N_{\mathcal{I}}}\left(t_{\phi}^{i}-t_{\theta}^{i}\right)^{2}\,,$ (14)
@@ 3.4 Relative Rotation from Correspondences
and $N_{\mathcal{I}}$ is the number of image correspondences.
@@
$f_{\mathcal{N}}^{(jk)}=\frac{1}{{N_{\mathcal{N}}}}\sum_{i}^{N_{\mathcal{N}}}\left\|{\bf N}_{j}[u_{j}^{i},v_{j}^{i}]-{\text{Norm}}(\mathrm{G}_{kj}{\bf N}_{k}[u_{k}^{i},v_{k}^{i}])\right\|^{2}\,,$ (17)
where Norm is the vector normalization operator and $N_{\mathcal{N}}$ the number of 3D correspondences.
@@
$f_{\mathcal{R}}^{(jk)}=\frac{1}{{N_{\mathcal{R}}}}\sum_{i}^{N_{\mathcal{R}}}\left\|{\bf N}_{j}[u_{j}^{i},v_{j}^{i}]-\Omega^{(jk)}({\bf N}_{k}[u_{k}^{i},v_{k}^{i}])\right\|^{2}\,,$ (18)
@@
Given 3 sets of $N_{\mathcal{I}}$ pixel correspondences, $N_{\mathcal{N}}$ 3D correspondences, and $N_{\mathcal{R}}$ reflection correspondences, we can optimize $f$ in Eq. 13 for the three Euler angles $\theta$, $\phi$, and $\eta$, and the two sets of parameters for the GBR transformations for both images $\mu_{1}$, $\nu_{1}$, $\lambda_{1}$, $\mu_{2}$, $\nu_{2}$, $\lambda_{2}$, under the constraints that $\lambda_{1}$ and $\lambda_{2}$ are positive.
Naive optimization with all the correspondences can easily be affected by erroneous correspondences which are unavoidable. We also empirically find that it is slow and susceptible to local minima especially due to the nonlinear operation $\omega_{\mathrm{r}}(\cdot)$. We instead derive a RANSAC-based, two-step algorithm that first estimates the combined transformation $\mathrm{G_{21}}$ using pixel and 3D correspondences and then decomposes it using reflection correspondences.
In the first step, given the pixel and the 3D correspondences, our algorithm recovers the combined transformation $\mathrm{G}_{21}\equiv\mathrm{G_{1}}^{-\mathrm{T}}\mathrm{R_{21}}\mathrm{G_{2}}^{\mathrm{T}}$ based on RANSAC [9]. For each of the $L_{1}$ iterations, we build a set consisting of randomly sampled $M_{1}$ pixels and $M_{1}$ 3D correspondences. In practice, we set $M_{1}$ to 4. We empirically find this to be the smallest necessary number of pixel and 3D correspondences. Please see the supplementary material for details. We hope to derive a theoretical justification of this number in our future work. For each set, we obtain an estimate of $\mathrm{G}_{21}$ by minimizing an objective function $f_{1}$ using an off-the-shelf solver for nonlinear optimization [31, 5]. $f_{1}$ is similar to the sum of $f_{\mathcal{I}}$, $f_{\mathcal{N}}^{(12)}$, and $f_{\mathcal{N}}^{(21)}$ in Eq. 13, though we compute $f$ using only the sampled correspondences. We then select a good estimate $\hat{\mathrm{G}}_{21}$ from them according to the number of pixels and 3D correspondences that are consistent with the estimate.
In the second step, given the estimate $\hat{\mathrm{G}}_{21}$ and the reflection correspondences, we obtain a combination of $\mathrm{G}_{1}$, $\mathrm{G}_{2}$, and $\mathrm{R}_{21}$ that is consistent with $\hat{\mathrm{G}}_{21}$ and most of the reflection correspondences. The key idea here is that, if we determine $\eta$, one of the three Euler angles of $\mathrm{R}_{21}$ that remains ambiguous, we can decompose $\hat{\mathrm{G}}_{21}$ into $\mathrm{G}_{1}$, $\mathrm{G}_{2}$, and $\mathrm{R}_{21}$ uniquely and analytically. Based on this, for each possible $\eta$, we compute the corresponding $\mathrm{G}_{1}$, $\mathrm{G}_{2}$, and $\mathrm{R}_{21}$, and select a solution that maximizes the number of reflection correspondences that are consistent with the decomposed transformations. The supplementary material provides more details and pseudo code.
3.4.1 Translation Estimation
Once $\mathrm{R_{21}}$ is recovered, by leveraging inlier pixel correspondences, we can also solve for the relative translation between the two views except for offsets regarding the two viewing directions [12]. That is, we estimate the translation vector $\mathbf{t_{21}}=(t_{x},t_{y},t_{z})$ using the constraint
where $r_{jk}$ is $(j,k)$ element of $\mathrm{R_{21}}$ and
$\tilde{u}^{i}=r_{11}u_{2}^{i}+r_{12}v_{2}^{i}-u_{1}^{i}\,,$ (21)
$\tilde{v}^{i}=r_{21}u_{2}^{i}+r_{22}v_{2}^{i}-v_{1}^{i}\,.$ (22)
Figure 4: Given two-view images of a textureless, non-Lambertian object, we first recover the surface normals and a reflectance map for each view using a single-view geometry reconstruction method [36]. We establish 3D and reflection correspondences with a novel deep feature extraction network to compute the relative camera pose from them.